OpenStudy (truwhovian11):

MEDAL AND FAN FOR THE FIRST PERSON TO HELP ME A new roller coaster has three seats in each row. The following shows how the number of seats, y, changes as the number of rows, x, changes. Which equation shows this relationship? y + 3 = x y = 3x 3y = x y = x + 3

OpenStudy (truwhovian11):

https://app151.studyisland.com/pics/251428q3.png

OpenStudy (thecalchater):

you see when x (rows equals 1) we have y =3 so which equation gives us that input/output?

OpenStudy (truwhovian11):

D? @TheCalcHater

OpenStudy (thecalchater):

nope because y=1+3 is 4 It would be B because y=1(3) is 3 and for every row we have 3 seats so for 10 rows we would have y=3(10) is 30 seats.

OpenStudy (truwhovian11):

oh ok thanks can you please help me with more

OpenStudy (lightlegendx):

Still need help? I'm Anaise, as you know :)

OpenStudy (truwhovian11):

what happend

OpenStudy (thecalchater):

Yeah I can help with a few more. (dont forget my fan+ medal :P)

OpenStudy (truwhovian11):

i always medal

OpenStudy (lightlegendx):

Nothing really

OpenStudy (thecalchater):

k close this one and open a new question and tag me.